A ball is thrown veticaly upward with an initial speed of 20m/s calculat 1)the time taken to return the acceleration. 2)the maximum night reach.

Grade:6

1 Answers

Dhrishi patel
5 years ago
Total time of flight =2u/g=2*20x10=4s
 
Max height reach=u square/2g=20m
Since dear student it was jst a formula for total time of flight i.e time of ascent +time of descent